OpenStudy (anonymous):

i was givin : y≤|x|-2 do you think it's suppose to be negative two or minus two?

OpenStudy (john_es):

You can split in two equations, \[x<0 \Rightarrow y\leq-x-2\]\[x\geq 0 \Rightarrow y\leq x-2\]

OpenStudy (anonymous):

but there is only suppose to be one, i have to graph it. is it normal to have a negative number right there, or do you think it is minus two?

OpenStudy (john_es):

You have to graph one, but it is needed to split in two, and you have to plot, truly, two parts of the same function due to the absolute value.

OpenStudy (john_es):

Also, the graph you will obtain is a region of the plane due to the inequality.

OpenStudy (anonymous):

 so is this correct?